I hate word problems! help please!
Number problem. If one-half of one integer is subtracted from three-fifths of the next consecutive integer, the difference is 3. What are the two integers?

Answer by checkley71(8403)

3/5(x+1)-x/2=3
(3x+3)/5-x/2=3 add the two fractions after finding the common denominator (10)
[2(3x+3)-5x]/10=3 now cross multiply
2(3x+3)-5x=3*10
6x+6-5x=30
x=30-6
x=24 answer.
24+1=25 answer.
proof
3/5*25-24/2=3
75/5-12=3
15-12=3
3=3